OpenStudy (anonymous):

Differentiate R(w) = csc(7w) using the Chain Rule. Can somebody answer and explain this for me? I'm really lost.

OpenStudy (anonymous):

The derivative of csc u is - csc u cot u du In your problem, let u = 7w. So its derivative is - csc (7w) cot (7w) times 7 Final answer: -7 csc(7w) cot (7w).

OpenStudy (phi):

csc(7w) = \( \frac{1}{\sin(7w)}\) you can make that to the -1 power: \[ \left( \sin(7w)\right)^{-1} \] now use the exponent rule: \( d x^n= n x^{n-1} dx \)